OpenStudy (zzr0ck3r):

prove that the sequence (-9/10)^n converges to 0

OpenStudy (anonymous):

I think you have to use the alternating series test.

OpenStudy (zzr0ck3r):

this is a sequence not a series

OpenStudy (zzr0ck3r):

no sumations

OpenStudy (zzr0ck3r):

it should have hte form \[|x_{n}-L| < epsilon\]

OpenStudy (zzr0ck3r):

bah equation editor is being lame

OpenStudy (anonymous):

oh, wait. yeah. sorry, i read it wrong. so you take the lim as n-> infinity of |r|^n. so lim n-> infinity of |9/10|^n = 0. so by the monotone convergence thm, it is convergent. also one of the properties of the limits of sequences is that if the limit of the absolute value of the sequence goes to 0, then so does the original series as well.

OpenStudy (zzr0ck3r):

I have to prove is using the form above

OpenStudy (anonymous):

hmm. maybe @TuringTest could help with that. I don't quite know how to put it in that form.

OpenStudy (anonymous):

did you end up getting it? :S

OpenStudy (dumbcow):

the way i would show it converges, is prove the limit as n->infinity is zero \[\lim_{n \rightarrow \infty}(-\frac{9}{10})^{n} = \lim_{n \rightarrow \infty}\frac{(-1)^{n}}{(\frac{10}{9})^{n}} = 0\]

OpenStudy (dumbcow):

sorry im not great at using epsilon delta proofs for these type of problems

OpenStudy (anonymous):

let \[\epsilon >0\] let \[N=\max( \lfloor \frac{-\ln(\epsilon)}{\ln(\frac{10}{9})} \rfloor+1;1)\] and let n > N then we will have \[\frac{-\ln( \epsilon)}{\ln(\frac{9}{10})} <N<n\] so , \[-\ln(\epsilon )<nln(\frac{10}{9})\] then \[\ln( \epsilon )>-nln(\frac{10}{9})=nln(\frac{9}{10})=\ln(\left|{(\frac{ -9}{10}})^n \right|)\] then wich mean that \[\epsilon>\left|{(\frac{ -9}{10}})^n \right|\] this finish the proof !

OpenStudy (zzr0ck3r):

ok thats exactly what I did, but didnt know if I was missing something tricky with ln. ty

OpenStudy (anonymous):

you're welcome ! :)

OpenStudy (zzr0ck3r):

so I dont know what the max thing is but I said let N > ln(1/ep)/ln(10/9) and n>N

OpenStudy (anonymous):

max is the greater number of both numbers ! and be careful ! you have to choose N as a positive integer !

OpenStudy (zzr0ck3r):

yeah, hmm I still dont understand why you used the max thing. We have not done anything like this so I just may need to think about it for a bit.

OpenStudy (anonymous):

just to make sure that my N is always a positive integer ! if epsilon >1 ! your N will be negative ! Mine will be positive =1

OpenStudy (zzr0ck3r):

so mine is not true? Im still a little iffy on whats going on with these problems. I learned this today...

OpenStudy (anonymous):

I meant by yours "ln(1/ep)/ln(10/9)" we are just dicussing ! so...Take it easy :)

OpenStudy (zzr0ck3r):

lol? Im easy. Im jsut asking cause im a little lost, and if the way Im doing it is not right im more lost:) Im not heated man. Dont know why it sounds that way but sorry:)

OpenStudy (zzr0ck3r):

I guess I dont understand why yet, but I just need to think about it more. What about the +1? what is that about?

OpenStudy (anonymous):

you know what's this \[\lfloor what's this \rfloor\] ?

OpenStudy (zzr0ck3r):

nope

OpenStudy (anonymous):

ohhh ! This is the reason of your lost !

OpenStudy (zzr0ck3r):

Im guessing thats only one of the reasons im lost:P

OpenStudy (anonymous):

trust me this the only one :) !

OpenStudy (zzr0ck3r):

so what is it?

OpenStudy (anonymous):

wait I looking for a link !

OpenStudy (zzr0ck3r):

ahh tyvm

OpenStudy (anonymous):

integer part means something to you ?!

OpenStudy (zzr0ck3r):

yeah

OpenStudy (anonymous):

that is the integer part ! :)

OpenStudy (anonymous):

now ! you agree that we have to choose N as a positive integer ! ?

OpenStudy (zzr0ck3r):

o that is notion for integer part, sweet. Yes tyvm

OpenStudy (zzr0ck3r):

notation*

OpenStudy (zzr0ck3r):

thats handy:)

OpenStudy (anonymous):

now ! did you understand !? :) !

OpenStudy (zzr0ck3r):

yes, forcing N to be positive.

OpenStudy (anonymous):

yeah !

OpenStudy (zzr0ck3r):

I love this stuff:) thankyou for your time.

OpenStudy (anonymous):

you're welcome ! just practice more ! you'll be fine !